Current flowing in conductors a and b is same. What is the ratio of the magnetic field produced around the conductor a at a distance of 5cm from the conductor to the magnetic field produced around the conductor b at a distance of 2cm from this conductor?


1
Expert's answer
2020-09-28T08:05:27-0400

"B_a=\mu\mu_0\frac{I}{2\pi a}"


"B_b=\mu\mu_0\frac{I}{2\pi b}"


"B_a\/B_b=b\/a=2\/5=0.4"